### v1

# 3D Anderson localization of light in disordered systems of dielectric particles

Abstract
We present the results of full wave numerical simulations of light transmission by layers of irregular dielectric particles that demonstrate 3D Anderson localization of light in disordered uncorrelated discrete media. Our simulations show that a high degree of disorder in a dense layer prevents transverse spreading of a propagating beam and allows anomalous conservation of its initial polarization state. A transition from the purely diffusive regime to a non-exponential time dependence when reaching the Ioffe-Regel criterion is observed in short pulse transmission measurements. The effect depends on the turbidity of the layer: increasing the volume fraction and the refractive index leads to the enhancement of the non-exponential behaviour caused by the disorder in the medium which is a clear signature of Anderson localization.
keywords:
disordered medium, light localization, light scattering
1 Introduction
Complete understanding and control of light propagation in disordered media can be advantageous in different applications in photonics like imaging and focusing in random media, random lasing, radiative transfer in particulate media and optical remote sensing [1, 2, 3, 4, 5]. However, this goal has not been achieved yet, in particular, for dense turbid media, due to the complexity of the problem for both analytical analysis and numerical simulations.
Light propagation in sparse media, in which scatterers are separated by distances larger than the wavelength, can be understood as a sequence of scattering events that are determined by the optical properties of individual particles, like single-scattering phase function and single-scattering albedo. If the number of such events is large enough, the optical transport process can be modelled using analytical multiple-scattering theories and the classic diffusion equation (e.g., [6, 7, 8]). In dense media with high concentration of scatterers, multiple scattering includes near-field interactions, and more complex transport phenomena may take place. For example, interference of counter-propagating waves may result in a weak localization and a coherent backscattering enhancement [9, 10]. Strong Anderson localization (AL) has been predicted for media with strong disorder and with multiple scattering not being suppressed by absorption [11, 12]. An important condition for the AL emergence is the sub-wavelength scale of coherent multiple scattering that can be expressed through the Ioffe-Regel criterion $kl^{*}<1$ for the mean free path length $l^{*}$ (where $k$ is the wavenumber) [13, 14]. This suggests a necessity of dense packing of particles in a disordered system in order to obtain AL. With high enough scattering strength of dielectric scatterers, that can be controlled by the variation of their refractive index, one can expect signatures of AL in the observations of light transmission by layers of particles. However, even non-absorbing layers with required properties and thickness $kL\gg 1$ can be quite opaque as it was already recognized by Anderson [11], which complicates the study of their transmission properties even further.
Generally, AL is characterized by a transition from diffusive propagation to a reduction and even complete halt of diffusion. An explanation involving trajectories with closed loops and increased return probability is often suggested [15, 16, 17]. However, the actual mechanism behind such dynamics has not been described yet in detail. The phenomenon is general in nature and can be observed for different kinds of waves, e.g., in acoustics [18], quantum systems [19, 20] and 2D photonic structures [21]. At the same time, the existence of 3D AL for electromagnetic waves in disordered uncorrelated media remains under debate. Experimental measurements of light transmission by layers of highly scattering particulate materials could shed light on the problem [17, 22, 23]. However, certain problems with precise control of the sample properties and accounting for the effects of absorption [24] and inelastic scattering [25, 26, 27] make this difficult. The most recent experimental results reported in [23] are debated because of ambiguous interpretation of the detected signature of AL that can be explained by fluorescence [27].
In theoretical studies of AL, one has to properly describe the disordered structure of a model medium and carefully take into account the multiple scattering near-field effects, including the polarization state of the scattered near field [28, 29, 16, 30]. Conventional multiple scattering theories cannot be applied here because of the small mean free path values at dense packing where the far-field single scattering properties of particles do not play role. In existing analytical models of the AL phenomenon, simplified descriptions of the media are used, e.g., a cloud of point scatterers [31] or a dielectric constant fluctuation around the average value $\epsilon(\vec{r})=\bar{\epsilon}+\delta\epsilon(\vec{r})$ [15]. If applied to the real ’white paint’-like 3D media, this can be insufficient.
Full wave numerical simulations can avoid approximations and unnecessary scattering effects considering pure interaction of the electromagnetic field with the target structure that may have arbitrary geometry. Modern numerical methods allow generation of particulate layer samples with high degrees of disorder, arbitrary shape of scatterers and packing densities [28, 32, 33]. At the same time, such an electromagnetic problem is multi-scale by definition. If the target object is much larger than the wavelength and has sub-wavelength structural features, accurate solutions require large computational meshes. This is a technical problem that can be resolved with the application of high-performance computing (HPC). However, only a few works on 3D numerical simulations and light localization in disordered and partially disordered structures have been reported [34, 35, 36, 37].
Observable characteristics of AL are the reduction of transverse spreading of the propagating light beam [38, 36, 34] and, in particular, time-delayed transmitted energy decay [39, 34]. These anomalies as unambiguous signatures of AL were found for the first time in a recent full wave numerical study of light transmission through 3D layers of densely packed perfect electric conductor (PEC) spheres [34]. That work suggests a possibility of the AL existence for electromagnetic waves in 3D in general. However, the case of dielectric non-absorbing spheres in layers with similar parameters and refractive indices up to $n=10$ showed only diffuse transport. Thus, the question whether AL can be observed for electromagnetic waves for disordered dielectric 3D structures remains an open problem to date.
In this work, we apply numerical simulations and HPC to study if conditions for AL can be created at $kl^{*}\lesssim 1$, according to the Ioffe-Regel criterion, and if AL signatures can be observed in light transmission by layers of dielectric particles. Reaching high enough degree of disorder at dense packing of irregular constituents, we observe a transition of classical diffusion characterised by the exponential time dependence of transmission $T(t)$ to AL with the $T(t)$ decay rate deviating from exponent. This behaviour is accompanied by the reduction of transverse spreading observed in the propagation of a focused beam. Thus, we demonstrate a possibility of AL in the ’white-paint’-like 3D systems of dielectric particles.
2 Numerical model
We apply the discontinuous Galerkin time-domain method (DGTD) [40, 32, 41, 42] to solve Maxwell’s equations for multi-particle layers illuminated by a plane wave as well as by a focused beam. For a qualitative benchmark of the DGTD light transmission computations for simplified structures we utilized the CST Microwave Studio software package, which is based on the Time Domain Finite Integration Technique (FIT) (see Supplementary material A.2).
We note that the $kl^{*}\lesssim 1$ criterion unconditionally requires dense packing of scatterers. Maximally random jammed packings of regular constituents, like spheres and regular polyhedra, always form hyper-uniform structures regardless of the packing algorithm [43, 44]. This kind of discrete media packed near their corresponding maxima, that depend on the type of constituents, have a reduced degree of disorder which influences their transmission properties (e.g., [45, 28] ). The problem can be solved by considering scatterers with random irregular shapes. To our knowledge, numerical simulations studying light localization in such systems have not been reported yet. Observation of AL in 3D layers of PEC spheres and its absence in the dielectric equivalent structures [34] provides us with an additional motivation to consider the case of irregular constituents.
Here, we simulate light transmission by mono-disperse layers with up to 20000 particles and use Gaussian random field shapes as constituents [28, 32, 46]. The Bullet physics engine [47] is applied to numerically pack them in systems with volume fractions up to $\rho=0.5$.
The constituent particle size is determined by the goal of satisfying the $kl^{*}\lesssim 1$ criterion and, possibly, to create conditions for light localization. The scale of free-space voids in a mono-disperse system becomes comparable to the particle size at large volume fractions. We take the particle size $kr=1.1$, where $r$ is the radius of the circumscribing sphere. In densely packed samples of compact irregular shapes their circumscribing spheres overlap making the distances between neighbour particles sufficiently small.
We consider two types of cylindric layers: thick ones for studying the focused beam propagation and thinner, but horizontally more extended ones for time-resolved transmission measurements. In the first case, the diameter and thickness are $kR=22.1$ and $kL=19.5$, respectively, and the number of particles is 20000 (Fig. 1a). In the second case, layers with various packing densities $\rho$ have larger diameter to thickness ratios and numbers of particles up to 19000. A layer sample with $kR=32$, $kL=14.5$ and $\rho=0.44$ and its vertical cross-section are shown in (Figs. 3ab). We note that the sample sizes are limited by our computational resources and the complexity of the unstructured mesh generation. For the DGTD simulations, we generate meshes with the particle systems surrounded by the perfectly matching layer (PML) to simulate an open boundary. We do not apply periodic boundary conditions in order to avoid possible uncertainties caused by the periodicity of the mesh. Cylindrical symmetry of samples may also affect the result. Therefore, we use samples with two types of the side boundaries: a randomly perturbed one and a smooth boundary. In the first case defects of the scale of $\sim\lambda$ that break such symmetry are introduced. In the second case a smooth boundary forms a dielectric discontinuity between the disordered medium and free space and, generally, one can expect whispering-gallery (WG) modes due to total internal reflection and cylindrical symmetry [48].
We consider a non-absorbing dielectric material and vary the refractive index $n$ in the range of $2.0\leq n\leq 3.0$.
3 Results and discussion
Figure 1: Propagation of a continuous, focused beam through a dense and a sparse layers of irregular particles. (a) and (c), samples with volume fractions of 0.48 (dense) and 0.16 (sparse), respectively. (b) and (d), steady-state near-field intensity distributions $|E|^{2}$ in XY, YZ and back-XZ planes for both layers. A white dotted line indicates a layer boundary with the diameter $kR=22.1$ and the thickness $kL=19.5$. The incident beam is $E_{x}$-polarized. The size of particles is $kr=1.1$, and the refractive index is $n=3.0$. The sparse structure shows a classic multiple-scattering pattern while dense packing strongly limits transverse spreading of the beam. An animation of a short pulse propagation, illustrating the transport mechanism in the dense case, is provided in the Supplementary material A.1.
Diffusive propagation of light is characterized by a transverse spread of energy and exponential energy decay in time as a result of multiple scattering. The geometry of the field spread experiences a change when particles are densely packed. This effect depends on the particle size and refractive index. Generally, at dense packing, far-field single scattering becomes impossible, and classic multiple scattering turns into percolation of light along random transmission channels [28, 29, 30]. Depending on the particle size and degree of disorder, this can lead to formation of a Caylee-tree-like channel [28] or a transversely localized propagation [38, 36]. This may occur even at a relatively small refractive index $n=1.5$ and $l^{*}\lesssim\lambda$ [28], which, however, should be insufficient for AL. Nevertheless, the diffuse field component appears to be significantly reduced under such conditions, which can be seen in the polarimetric measurements of backscattering from high-albedo powders with different packing densities [49, 50, 51]. Little free space between the neighboring scatterers in a medium reduces the role of diffraction after each scattering event and, consequently, the contribution of the diffuse field [28]. The measured polarimetric features produced by single particles on the surface, that are suppressed by unpolarized diffuse multiple scattering in sparse samples, get enhanced in compressed ones. Hence, in the search for real AL, we consider samples with parameters that are close to those predicted in the theoretical works.
In Fig. 1, we study focused beam propagation through a thick layer of particles with a size $kr=1.1$ and a refractive index $n=3.0$. The internal field patterns are compared for a dense and a sparse structures of $\rho=0.48$ (Fig. 1a) and 0.16 (Fig. 1c), respectively. A continuous $E_{x}$-polarized focused beam incident on the front side in each case propagates in $Z$ direction. Steady-state intensity distributions as cross-sections in different planes are shown in Figs. 1b and 1d. The ”XY” panels correspond to the cross-sections near the back sides of the samples.
In the sparse case (Fig. 1d), we observe classic scattering by particle groups and individual particles, which appear to be weak scatterers at large separations from each other. In a dense system, the size of voids is reduced to $kl^{*}\lesssim 1$, and a focused beam tends to preserve its diameter along the path, re-emitting transmitted energy in the center of the back side. Small isotropic spreading occurs due to the leakage of energy through the evanescent field coupling. Therefore, the intensity of the beam decreases, but its geometric cross-section remains constant at the distance of the layer thickness. The dynamics of the process can be seen in the animation in the Supplementary material A.1. It shows the evolution of a focused beam, percolating through a dense layer without transverse spreading. Light propagation evolves mainly through evanescent coupling between individual scatterers without conventional free-space transport of the scattered waves. This reminds us of an evolution of a cellular automata (CA) in which the vector field state in an elementary cell at each time step is determined by its nearest environment, i.e. the shape of neighbour particles or the shape of a sub-wavelength void that is formed by the neighbors. The laws of interaction of the vector field and the air-dielectric interface become the CA rules. Then, a sequence of such interactions determines the evolution of the entire vector field from the initial state, resulting in a forward pulse propagation.
The fundamental property of the observed transport mechanism is that most of the energy of the propagating beam is concentrated in the free-space cavities, and little is refracted into the dielectric particles. A wave can couple to the neighbor-particle interface while being still coupled to the previous one propagating in a void between two interfaces. On the other hand, these interfaces are formally disconnected and belong to different scatterers. Thus, the field appears to be constantly coupled and one can say that there are no more ”scattering events”. Therefore, the term ”free path length” becomes irrelevant in its classic definition (see also [14]). $kl^{*}<1$ implies also an extremely small mean free time $t^{*}$, and the near field, being coupled to a particle, most often reaches its neighbour at time $t\ll T$, where $T$ is the wave oscillation period. Thus, we refer to the $kl^{*}$ parameter only to characterise the size of the voids in a particulate medium with respect to the wavelength.
Furthermore, we note that the wave packet is partially preserved along the propagation path. Its effective wavelength becomes smaller, $\lambda_{eff}\approx 0.6\lambda_{inc}$, increasing the effective size of particles to $k_{eff}r\approx 1.8$. Nevertheless, the condition $kl^{*}\lesssim 1$ is preserved for most of the voids.
Figure 2: Propagation of a continuous pulse through the dense layer shown in Fig. 1a. Steady state distribution of the components of the electric near field $\vec{E}$ (XZ-plane) show that the transmitted pulse is not totally depolarized but, on the contrary, preserves most of its initial polarization state $E_{x}$ after transmission. The $E_{y}$ component as well as $E_{z}$ is excited locally in the voids between the particles along the propagation path but is not transmitted directionally.
Diffusive propagation in a disordered medium depolarizes incident radiation after many scattering events that are strongly stochastic. A steady-state distribution of the electric field component $E_{x}$ in Fig. 2 computed for a dense layer shows a counter-intuitive conservation of the original polarization state of the propagating beam at the exit. The $E_{y}$ and $E_{z}$ components are excited along the path and play a role in the field coupling process but they do not propagate directionally. This fact can explain the halt of transverse spreading: the propagating beam ”prefers” cavities where the initial polarization direction can be preserved and, hence, the $\vec{k}$-vector bending is avoided and the direction of propagation is not changed.
Figure 3: (a) A layer of 19000 irregular particles packed with a volume fraction of 0.44 and with dimensions of $kR=32$ and $kL=14.5$. The size of particles is $kr=1.1$ and the refractive index is $n=3.0$. (b) vertical cross-section of the layer, showing intentional defects on the side boundary and the characteristic scale of voids $kl\lesssim 1$. (c) Normalized transmission of a short pulse by layers with volume fractions from $\rho=0.17$ to $0.5$ as a function of time $T(t)$, exhibiting a transition from the exponential regime to a non-exponential decay at longer times. (d) Time-dependent diffusion coefficient $D(t)$ obtained by local exponential fitting of $T(t)$. Both axes have logarithmic scale. A dense layer with $\rho=0.44$ demonstrates a possible transition to $t^{-1}$ (dashed line) dependence at times $t>20t_{D}$.
Figure 4: Simulation results for layers with different volume fractions and smooth side boundaries. The sample dimensions are of $kR=26$ and $kL=12$, respectively. The size of particles is $kr=1$ and the refractive index is $n=3.0$. (a) Normalized transmission of a short pulse by layers with volume fractions from $\rho=0.1$ to 0.5 as a function of time $T(t)$, exhibiting a transition from the purely exponential regime to a non-exponential decay at $\rho\gtrsim 0.3$. (b) Time-dependent diffusion coefficient $D(t)$ obtained by local exponential fitting of $T(t)$. Dense layers demonstrate a $t^{-1}$ (dashed lines) dependence at times $t>20t_{D}$. (c) and (d), Snapshots of the internal electric field (XY-cross-sections, $E_{x}$ component) distributions showing a whispering-gallery wave propagating along the boundary at time $t=25t_{D}$ for a dense ($\rho=0.44$) and a moderately sparse ($\rho=0.27$) layers. In the last case $T(t)$ is exponential despite the whispering-gallery effect.
Transverse localization in a dense structure may be accompanied by the changes in the time dependence of transmission, analogous to those observed for dense packings of PEC spheres in [34]. Here we perform a related analysis. We simulate time-resolved transmission by layers $T(t)$ varying their turbidity, i.e. the volume fraction and the refractive index. At first, we consider layers with perturbed side boundaries and with dimensions of $kR=32$ and $kL=14.5$. The volume fraction $\rho$ is varied between 0.17 and 0.5.
[…]
Non-exponential decay can be formally represented by means of a time-dependent diffusion coefficient $D(t)$. At the localized state, it decreases as $t^{-1}$ according to the solution of the diffusion equation for the case of $kl^{*}<1$ in time domain [39]. Figure 3d shows time resolved diffusion coefficients $D(t)$ deduced from the computed $T(t)$ by local exponential fitting. One can see a transition from small variation of $D(t)$ in sparse structures to a steep function at $\rho>0.44$ and times $t>15t_{D}$.
Evidently we would have to extend the observation time limit in order to study the evolution of $T(t)$ at longer times and to find out if $D(t)$ can reach the predicted $t^{-1}$ dependence. This could be achieved by increasing sample thickness and diameter roughly by a factor of two or more. This would correspond to $kL\gtrsim 30$ and a number of particles $N>10^{5}$ which makes the problem computationally extremely expensive, even for modern large scale HPC clusters. Another way to hinder energy decay in time is to create conditions for a whispering-gallery effect at the side boundary of a cylindric layer. For this purpose we did computations for layers with smooth boundaries and a particle size of $kr=1$. The results for $T(t)$ and $D(t)$ dependencies at different volume fractions from $0.1$ to $0.5$ are shown in Figures 4ab. The dimensions of the layers are $kR=26$ and $kL=12$. A densely packed sample with $\rho=0.44$ is shown in the inset in Figure 4a. Figures 4cd represent snapshots of the electric field distributions ($E_{x}$ field component) in the horizontal XY-plane at time $t=25t_{D}$ for a dense ($\rho=0.44$) and a moderately sparse ($\rho=0.27$) layers. The WG waves, excited by the incident pulse due to the smooth boundaries of samples with cylindric symmetry, are coupled to the side boundary due to total internal reflection and allow us reaching $t\approx 40t_{D}$ for relatively small samples. At the same time the field distributions remain random in the bulk. Despite the WG effect, present in both cases, the structures show different transmission properties. The $T(t)$ plot in Figure 4a clearly shows a transition from purely exponential decay to localization at large volume fractions and times $t>20t_{D}$. This occurs at $\rho\gtrsim 0.3$. The WG effect is observed in all the layers above and below this value, even in the sparse structures for which the the condition $l^{*}\lesssim\lambda$ holds. This proves that it is not the WG mode itself but the particle packing density which is causing the non-exponential behaviour of $T(t)$. The diffusion coefficient $D(t)$ at times $t\gtrsim 20t_{D}$ obtained for the dense structures with $\rho\gtrsim 0.44$ follows a $t^{-1}$ fit in Figure 4b demonstrating that we have reached the AL regime. We note that this result is consistent, in terms of time and the volume fraction threshold, with that obtained for layers of PEC spheres in [34].
Figure 5: (a) Normalized transmission of a short pulse through a layer with a volume fraction $\rho=0.44$ and a smooth boundary as a function of time $T(t)$ at different refractive indices. (b) Time-dependent diffusion coefficient $D(t)$ as obtained by a local exponential fitting of $T(t)$. A localization signature $D(t)\sim t^{-1}$ emerges at times $t>20t_{D}$ and $n=3.0$.
[…]
Figure 6: A snapshot of the near field $|E|^{2}$ excited by the propagation of a short pulse through a dense layer with $\rho=0.48$ and $n=3.0$ shown in Fig. 1a along $Z$ axis at times $t=2t_{D}$ (a) and $t=20t_{D}$ (b). A fraction of light is trapped by the disordered air-dielectric interface through evanescent coupling. This results in a constant flow along complex, highly disordered paths which slows down the energy decay through the layer boundaries. An animation showing the time dynamics of the near field can be found in the Supplementary material A.1.
Our results imply a mechanism that slows down the energy decay at dense packing. Apparently, it should be connected with the percolation-like light transport in a thick, dense layer we analyzed above. Visualization of the field dynamics excited by a short pulse yields a clue to this mechanism. One can conventionally split up the propagating field into two components. The first one is the wave packet of the incident pulse that quickly diffuses through a layer. The second one is created by the energy leaked from the packet through the evanescent field-coupling. Being coupled once, this component remains trapped by the air-dielectric interface, which, together with the free-space cavities, represents a geometrically complex 3D structure. Despite little free space at dense packing, such a structure possesses a long-range connectivity, i.e. arbitrary points in a void can be connected with another point in an arbitrary free-space location. The paths connecting such arbitrary points can be much longer than the sample dimensions.
In Fig. 6, one can see snapshots of the near-field intensities inside the thick layer shown in Fig. 1a at a time right after the pulse passage $t=2t_{D}$ and at the longer time $t=20t_{D}$. An animated field evolution for a period of time after $2t_{D}$ can be found in the Supplementary material A.1. The trapped field uniformly fills the sample volume and propagates along complex paths, constantly changing direction and forming hotspots in the voids. The paths can be not necessarily closed loops, but they are apparently sufficiently long. The field decay occurs only when the field reaches the sample boundaries. Thus, the energy in the interior needs substantial time in order to reach the boundary, much longer than the time of propagation along a straight line or a sequence of straight lines between scattering events in a sparse medium. The character of such an internal field evolution does not significantly change at longer times, and the decay of hotspots is not monotonous in each location. They demonstrate different behaviour in such a disordered system: those that are excited directly by the transmitted pulse decay while others arise in new locations. The decay rate is also different for different hotspots, i.e. they may have different life times. Some are decaying within $\approx 5$ oscillation periods while others are much more stable. Non-monotonous decay and enhancement with time may also take place in the same location, which proves a high geometrical complexity of the paths. If the energy of a path is leaked at the sample boundary, it is lost in the total decay. A leakage in the interior transfers energy to new paths with similar spatial scales and degrees of complexity, forming a complex percolation graph. Thus, the entire process appears to be a sophisticated spatiotemporal 3D flow of energy and should be a subject for further studies. In particular, such a mechanism can be relevant for low-absorbing materials as it involves field-interface interaction and propagation in free space. The role of the refractive index also becomes clear in this situation. It determines the balance between the energy transmitted by particles and chains of particles, enhancing a diffusive process, and the energy flowing through a random system of connected cavities with long path lengths that participates in localization.
4 Conclusions
We observe clear signatures of AL in the results of the full wave numerical simulations of light transmission by 3D dielectric particulate layers with high degrees of structural disorder. A transition from diffusive transport to localization occurs at approaching the Ioffe-Regel criterion for the mean free path length $kl^{*}\lesssim 1$ in densely packed systems and at the refractive index of $n=3.0$, which is consistent with the predictions of theoretical models. Simulation of the focused beam propagation in a dense layer demonstrates an absence of transverse spreading and anomalous conservation of the beam’s linear polarization state in contrast to the classical light scattering in a sparse analog structure. Time-resolved transmission of a short pulse in layers with similar densities and constituent particle sizes exhibits a non-exponential energy decay already at times $t>15t_{D}$. For observations at $t\gg t_{D}$ larger structures must be considered which makes the problem computationally very expensive even using HPC. Alternatively, we extend the observation time retaining sample cylindric symmetry. This allows excitation of whispering gallery modes that hinder field decay from a sample. Therewith we found that the time-resolved diffusion coefficient varies in agreement with the self-consistent theory of localization at times $t>20t_{D}$ and volume fractions $\rho\gtrsim 0.44$. Thus, Anderson localization of light is possible in dense highly disordered 3D dielectric structures which motivates a continued experimental search of the AL characteristics in uncorrelated disordered media, with realistic material refractive indices not exceeding $n=3.0$.
[…]
Methods We performed numerical simulations using a self-developed light scattering code based on the open source DGTD Maxwell solver MIDG published by Tim Warburton [52]. The CST Microwave Studio software package based on the Time Domain Finite Integration Technique was used for a qualitative benchmarking the DGTD simulation results for simplified setups. In our code we implemented plane-wave and focused beam lights sources, boundary conditions and material description. Here, we use continuous and Gaussian pulses. A plane wave source is realized using the total field/scattered field technique. For simulations with a focused beam we implemented the plane wave superposition method [53]. An open computational domain is pretended with the PML absorbing boundary condition. For the time-resolved transmission measurements we illuminated target layers with a linearly polarized short plane wave pulse. The intensity of the transmitted signal is recorded, correspondingly, in a monitor plane behind the layer. The intensity curves in all plots are normalized by the peak intensity of the transmitted pulse. For generation of the particulate samples with different dimensions and controlled filling fractions we use the Bullet Physics engine [47]. This is an open-source C++ library that can be also used as an add-on in the Blender software. It allows simulation of dynamics and collision of multiple arbitrary 3D shapes in time domain. Thus, simulating natural powder mechanics samples with realistic topologies can be created. For sparse distributions of the pre-generated random irregular shapes we simulated their free fall on a substrate in a closed cylindrical volume. The sets of particles are created with a self-developed shape generator based on the Gaussian random field approach [46]. 3D models of the generated layers were then used for tetrahedral mesh generation. The spatial resolution in the regions with dielectric material with $n=3.0$ was $\sim 15$ tetrahedral cells per central wavelength of the incident pulse and the nodal expansion order in the DGTD numerical scheme was $N=3$. All the DGTD simulations were done on the HPC cluster Noctua 2 of Paderborn University using up to 30 cluster nodes per simulation.
Appendix A Supplementary material
A.1 Near field dynamics
Animated evolution of the near field in time helps to better represent the field dynamics in such a complex structure as a discrete random medium consisting of densely packed irregular particles. To support our analysis of the static field snapshots in Figure 1b we additionally simulate propagation of a short focused pulse in the same dense layer sample (Figure 1a) with the same model parameters. Animation A.1 represents an $XZ$ cross-section of the computational domain with the near-field intensity distribution and shows a cellular-automata-like behavior of the local field hotspots in the sub-wavelength voids between the closely neighboring particles.
Figure A.1: Animated propagation of a short focused pulse in a dense layer shown in Figure 1a.
Animation A.2 demonstrates dynamics of the near field that is trapped by the air-dielectric interface of the constituent particles after propagation of a short plane-wave pulse in a short period near $t=2t_{D}$. At high degree of disorder and in the absence of large voids the evanescent coupling mechanism slows down the energy decay from the system.
Figure A.2: Animated dynamics of the near field in a dense layer shown in Figure 1a excited by a short plane wave pulse.
A.2 Pulse transmission by layer samples with reduced dimensions
Here we simulate short-pulse transmission by layers with reduced dimensions in order to study the transition to AL increasing the geometrical complexity through the layer thickness and the number of particles. In addition we do test simulations using the CST Microwave Studio software package. It is based on the Time Domain Finite Integration Technique, so, one can qualitatively benchmark our DGTD results with a different numerical method and a different electromagnetic solver.
We take a cylindrically symmetric sample with $kR=26$, $kL=12$ and $\rho=0.44$ as a reference and consider reduced structures with thicknesses $kL=6$ and 3 and an elementary bi-layer structure. The refractive index is fixed at $n=3.0$. The results of the simulated transmission $T(t)$ for four layers are presented in Figure A.3a. A bi-layer shows purely exponential decay even though a WG mode is excited. The case of $kL=3$ is equivalent to a doubling the thickness of the bi-layer sample. This appears to be enough to make the $T(t)$ curve non-exponential. In all four cases a WG mode is excited allowing us reaching the $25-40t_{D}$ region even with such a small thickness. Further thickness increasing leads to stronger deviations of exponent and faster reduction of the diffusion coefficient $D(t)$ (Figure A.3b). However, only with $kL=12$ we obtain a qualitative change at longer times and reach the $t^{-1}$ fit.
Figure A.3: (a) Normalized transmission of a short pulse through layers with cylindric symmetry and different thickness $kL$ as a function of time $T(t)$. The diameter and volume fraction are $kR=26$ and $\rho=0.44$, correspondingly, and the refractive index is $n=3.0$. The energy decay from a simple double layer of particles is exponential while thicker layers demonstrate increasingly non-exponential transmission. (b) Diffusion coefficient $D(t)$ obtained by local exponential fitting of $T(t)$. $D(t)$ quickly evolves from a constant to a $t^{-1}$ fit (dashed line) with increasing thickness. (c) Vertical cross-sections of the samples with different thickness.
Figure A.4: Normalized transmission of a short pulse through a small layer of 2000 particles packed with volume fraction $\rho=0.5$ (inset) and with dimensions $kR=15$ and $kL=3$ as a function of time $T(t)$ at different refractive indices. Simulations are done with the CST Microwave Studio software package.
For the computations with the CST FIT solver we take a cylindric sample with thickness $kL=3$ (a four-particle thick layer) and reduced diameter $kR=15$. Due to parallelization restrictions of CST MWS we need to simplify the target system. With the particle size $kr=1$ and the volume fraction $\rho=0.5$ the number of particles becomes relatively small in this case, N=2000. With fixed sample parameters we vary the material refractive index $n$ from 2.0 to 3.0. For the CST computations we use a polyhedral description of the model sample and, then, a hexahedral mesh is generated by the software internally. The electromagnetic problem is solved in time domain simulating propagation of the short plane-wave pulse, similarly to our DGTD method setup.
The result in Figure A.4 is qualitatively similar to the $T(t)$ refractive index dependence in Figure 5a obtained for a much larger system with the DGTD method.
Summarizing all our results we note that increasing three parameters: layer thickness, volume fraction and refractive index act in very similar ways causing transition from an exponential dependence of $T(t)$ to a non-exponential behavior with qualitative change at longer times.